# A Multiobjective Mathematical Model for Optimal Irrigation Water Allocation

### v1

@@ 7 Conclusion
This study contributes to the growing efforts in sustainable water management by introducing a practical optimization framework designed to strike a balance between economic gain and environmental responsibility. By using three widely used techniques, gamultiobjectivite,Paretosearch and scalarization, the model provides an accessible and effective tool that can support planners in making more informed and thoughtful decisions when managing scarce water resources. The findings suggest that smart allocation strategies can lead to noticeably higher economic benefits while also safeguarding environmental flows. The results show that scalarization, in particular, is well-suited to handle trade-offs between conflicting goals. The model can also successfully consider real-world operational constraints, making it relevant and applicable in practice. Moreover, since the analysis is focused on a single irrigation project in a tropical monsoon area, the findings are also applicable in arid and semi-arid areas.
The study has some limitations on the type of data used. It follows a deterministic model where data are certain and previously known. It doesn’t fully reflect uncertainties or covers Stochastic or fuzzy aspects like changing climate conditions, fluctuations in crop yields, or shifting policy and social dynamics.
There is a scope to enhance the model by including different area boundaries for each crop on its demand, canal capacity constraints and crop scheduling based on land and crop types, as well as by adding deep percolation, soil moisture, dam storage, overflow, and reservoir evaporation constraints in future work. In addition, connecting the model with socio-economic and policy frameworks could help to better understand its broader impact and make it even more useful for decision-makers at both local and national levels.
Acknowledgment: We acknowledge Chatgpt(5) AI tools, which are used to rephrase, edit and polish the author’s written text for spelling, grammar, or general style.
@@ 8 Funding and/or Conflicts of interests/Competing interests
The authors declare that no funds, grants, or other support was received during the preparation of this manuscript. The authors have no relevant financial or non-financial interests to disclose.

### v2

@@ 8 Conclusion
This study contributes to the growing efforts in sustainable water management by introducing a practical optimization framework designed to strike a balance between economic gain and environmental responsibility. By applying three widely used multiobjective optimization approaches, implemented through the MATLAB functions ‘gamultiobj’ and ‘paretosearch’ alongside a scalarization method, the framework provides a transparent means of exploring trade-offs in irrigation water allocation under scarcity. The findings suggest that smart allocation strategies can lead to noticeably higher economic benefits while also safeguarding environmental flows. The results show that scalarization, in particular, is well suited to handle trade-offs between conflicting goals. The model can also successfully consider real-world operational constraints, making it relevant and applicable in practice. Moreover, since the analysis is focused on a single irrigation project in a tropical monsoon area, the findings are also applicable in arid and semi arid areas. The primary contribution of this study lies in methodological comparison and optimization performance assessment rather than direct operational prescription.
The study has several limitations related to data and modeling assumptions. It adopts a deterministic framework in which input data are assumed to be known with certainty and relies solely on dry year conditions. As a result, the model does not capture uncertainties associated with climate variability, crop yield fluctuations, or changes in target environmental flow requirements. Moreover, fuzzy and socio-policy dynamics are not explicitly represented.
There is substantial scope to further enhance the proposed modeling framework by incorporating additional operational and hydrological complexities. Future work may include crop specific area constraints based on demand, canal capacity limits, and crop scheduling decisions linked to land suitability and crop types. The model can also be extended to represent key hydrological processes such as deep percolation, soil moisture dynamics, dam storage, overflow, and reservoir evaporation, particularly during periods of excess rainfall. Introducing these elements would improve the representation of water availability and storage behavior under monsoon conditions. In addition, allowing the target Environmental Flow to vary seasonally would better reflect natural hydrological variability and improve ecological realism. Such seasonally adaptive targets would enable a more accurate assessment of trade-offs between irrigation withdrawals and downstream environmental protection across wet and dry periods. Finally, integrating the framework with socioeconomic and policy modules would enhance its relevance for decision makers at both local and national levels.
Acknowledgment: We acknowledge ChatGPT (5) AI tools, which are used to rephrase, edit, and polish the author’s written text for spelling, grammar, or general style.
@@ 9 Funding and/or Conflicts of interests/Competing interests
The authors declare that no funds, grants, or other support were received during the preparation of this manuscript. The authors have no relevant financial or non-financial interests to disclose.